# Accuracy Tradeoff for Out-of-Distribution Few-shot Learning

### v1

@@ 1 Introduction
There is extensive interest from the computer vision community for training classifiers that are robust to distribution shifts. Pioneering works in this area [23, 65, 46] focused on optimizing for simple shifts in the image distribution, such as sketch-to-real adaptation. As the topic evolved, the community proposed increasingly harder adaptation problems by eliminating some restrictive assumptions. For the domain generalization (DG) problem [66], we do not assume access to unlabeled target data; for the cross-dataset generalization (XD) problem [68], we allow source and target label spaces to be different; and for the parameter efficient learning (PEFT) problem [20, 55], we impose a tight budget on the number of parameters that can be tuned. Our work lies at the confluence of these three topics. Similar to CoOp [68] and MaPLe [25], we do assume access to labeled few-shot generic source data, such as ImageNet. Since we assume nothing about the relationship between source and target datasets, this setting can be more useful in practice than strict zero-shot learning. In this paper, we propose two parameter efficient few-shot methods, called word and descriptor soups, that finetune vision-language (VL) models to generalize to target datasets which may contain unseen labels and/or shifts in the image distribution. Our methods achieve state-of-the-art on some benchmarks without additional gradient-based tuning, but can also improve state-of-the-art gradient-based finetuning methods with an additional diversity loss.
@@ 3.4 Diversity loss
CoOp [68$\dagger$ 71.5 63.88 59.3
Co-CoOp [67$\dagger$ 71.0 65.74 59.9
@@ Discussion
An ablation study on our soup methods with varying $m$ is presented in Fig. 5. On both benchmarks, our word soup performs best for all $m$. We note that the word soup with $m=2$ already outperforms all ZS baselines for all values of $m$ up to 64. This result indicates that, unlike state-of-the-art ZS methods, ensembling is not the main ingredient of our method.
Parameter Efficiency
Computational Efficiency
@@ Acknowledgements
# @ means matrix multiplication in Python.
# ’+’ means concatenation when operating on lists.
# Inputs: L2-normalized image_embeddings, y_truth
# classnames: list of classnames in English
# model: CLIP-style model
# descriptions: list of descriptions from an LLM
# Hyperparameters: m (number of members in the soup)
descriptions = [’which has legs.’, which can swim.’,  ]
[…]
y_truth):
scores = image_embeddings @ text_embeddings.T
return (scores.argmax(dim=1) == y_truth).mean()
[…]
d = tokenizer([’a photo of  + classname + ’,  +
%**** main.tex Line 750 ****description for classname in classnames])
return normalize(model.encode_text(d))
[…]
for description in descriptions:
text_embeddings = get_description_embeddings(
description)
accuracies.append(get_accuracy(
image_embeddings, text_embeddings, y_truth))
# sort descriptions by accuracies
descriptions_sorted = descriptions[
accuracies.sort(descending=True).indices]
[…]
soup, accuracy = [descriptions_sorted[0]], accuracies[0]
[…]
for description in descriptions_sorted:
soup_embeddings = stack(
[get_description_embeddings(description)
for description in soup + [description] ] )
text_embeddings = normalize(
soup_embeddings.mean(dim=0))
if get_accuracy(image_embeddings,
text_embeddings, y_truth) > current_acc:
%**** main.tex Line 775 ****soup = soup + [description]
[…]
@@
# Hyperparameters: k0, k1, m, patience
# Inputs: L2-normalized image_embeddings, y_truth
[…]
accuracies = []
for word in words:
text_embeddings = get_description_embeddings(word)
accuracies.append(get_accuracy(
image_embeddings, text_embeddings, y_truth))
# sort descriptions by accuracies
words = words[accuracies.sort(descending=True).indices]
[…]
for repeat m times:
first_word = random.shuffled(words[0:k0])[0]
%**** main.tex Line 800 ****word_chain = first_word
accuracy = get_accuracy(image_embeddings,
get_description_embeddings(word_chain), y_truth)
words_k1 = random.shuffled(words[0:k1])[0:patience]
[…]
for word in words_k1:
text_embeddings = get_description_embeddings(
word_chain +   + word)
next_accuracy = get_accuracy(image_embeddings,
text_embeddings, y_truth)
if next_accuracy > accuracy:
word_chain = word_chain +   + word
soup = soup + [word_chain]
[…]
@@ Appendix C Token offset trick (for Descriptor Soup)
CoOp [68$\dagger$ 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88 64.20 47.99 49.71 75.21 59.3
Co-CoOp [67$\dagger$ 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74 64.07 48.75 50.63 76.18 59.9

### v2

@@ 1 Introduction
There is extensive interest from the computer vision community for training classifiers that are robust to distribution shifts. Pioneering works in this area [23, 67, 47] focused on optimizing for simple shifts in the image distribution, such as sketch-to-real adaptation. As the topic evolved, the community proposed increasingly harder adaptation problems by eliminating some restrictive assumptions. For the domain generalization (DG) problem [68, 55], we do not assume access to unlabeled target data; for the cross-dataset generalization (XD) problem [70], we allow source and target label spaces to be different; and for the parameter efficient learning (PEFT) problem [20, 57, 37], we impose a tight budget on the number of parameters that can be tuned. Our work lies at the confluence of these three topics. Similar to CoOp [70] and MaPLe [25], we do assume access to labeled few-shot generic source data, such as ImageNet. Since we assume nothing about the relationship between source and target datasets, this setting can be more useful in practice than strict zero-shot learning. In this paper, we propose two parameter efficient few-shot methods, called word and descriptor soups, that finetune vision-language (VL) models to generalize to target datasets which may contain unseen labels and/or shifts in the image distribution. Our methods achieve state-of-the-art on some benchmarks without additional gradient-based tuning, but can also improve state-of-the-art gradient-based finetuning methods with an additional diversity loss.
@@ 3.4 Diversity loss
CoOp [70$\dagger$ 71.5 63.88 59.3
Co-CoOp [69$\dagger$ 71.0 65.74 59.9
@@ Discussion
An ablation study on our soup methods with varying $m$ is presented in Fig. 5. On both benchmarks, our word soup performs best for all $m$. We note that the word soup with $m=2$ already outperforms all ZS baselines for all values of $m$ up to 64. This result indicates that, unlike state-of-the-art ZS methods, ensembling is not the main ingredient of our method. Additional ablation studies are presented in Appendix E.
Parameter Efficiency and Computational Efficiency
A discussion regarding efficiency of our methods is deferred to Appendix E.
@@ Acknowledgements
{minted}
[fontsize=]python # ’@’ means matrix multiplication in Python. # ’+’ means concatenation when operating on lists. # Inputs: L2-normalized image_embeddings, y_truth # classnames: list of classnames in English # model: CLIP-style model # descriptions: list of descriptions from an LLM # Hyperparameters: m (number of members in the soup) descriptions = [’which has legs.’, ’which can swim.’, … ]
def get_accuracy(image_embeddings, text_embeddings, y_truth): scores = image_embeddings @ text_embeddings.T return (scores.argmax(dim=1) == y_truth).mean()
def get_description_embeddings(description): d = tokenizer([’a photo of ’ + classname + ’, ’ + description for classname in classnames]) return normalize(model.encode_text(d))
accuracies = [] for description in descriptions: text_embeddings = get_description_embeddings( description) accuracies.append(get_accuracy( image_embeddings, text_embeddings, y_truth)) # sort descriptions by accuracies descriptions_sorted = descriptions[ accuracies.sort(descending=True).indices]
# initialize with best descriptor soup, accuracy = [descriptions_sorted[0]], accuracies[0]
# greedy selection for description in descriptions_sorted: soup_embeddings = stack( [get_description_embeddings(description) for description in soup + [description] ] ) text_embeddings = normalize( soup_embeddings.mean(dim=0)) if get_accuracy(image_embeddings, text_embeddings, y_truth) > current_acc: soup = soup + [description]
return soup[:m]
@@
{minted}
[fontsize=]python # Hyperparameters: k0, k1, m, patience # Inputs: L2-normalized image_embeddings, y_truth
words = ["the", "of", "and", … ] accuracies = [] for word in words: text_embeddings = get_description_embeddings(word) accuracies.append(get_accuracy( image_embeddings, text_embeddings, y_truth)) # sort descriptions by accuracies words = words[accuracies.sort(descending=True).indices]
soup = [] for repeat m times: first_word = random.shuffled(words[0:k0])[0] word_chain = first_word accuracy = get_accuracy(image_embeddings, get_description_embeddings(word_chain), y_truth) words_k1 = random.shuffled(words[0:k1])[0:patience]
# greedy selection for word in words_k1: text_embeddings = get_description_embeddings( word_chain + " " + word) next_accuracy = get_accuracy(image_embeddings, text_embeddings, y_truth) if next_accuracy > accuracy: word_chain = word_chain + " " + word soup = soup + [word_chain]
return soup
Limitations
Similar to many related works, the main limitation of our work is that we require the source dataset to cover a broad range of classes (e.g. ImageNet). As a counter example, we cannot hope to train on pets classification and generalize to ImageNet. We highlighted this limitation in Table 1 of the main paper (top) with qualitative examples.
@@ Appendix B Additional Word Soup Motivation
INet
Caltech
Pets
Cars
Flowers
Food
Aircraft
SUN
DTD
EuroSAT
UCF
Mean
INet-V2
Sketch
INet-A
INet-R
Mean
CLIP ZS 67.1 93.3 89.0 65.4 71.0 85.7 25.0 63.2 43.6 46.7 67.4 65.02 61.0 46.6 47.2 74.1 57.22
Word soup 68.8 94.1 89.5 65.9 72.6 86.3 26.1 67.2 45.3 53.9 67.8 66.87 62.6 49.0 50.4 77.0 59.73
Vanilla CoOp 68.7 94.4 90.2 66.1 70.9 85.8 26.0 66.7 47.4 50.1 68.9 66.63 61.9 48.6 49.8 76.7 59.26
    + Word soup 69.1 94.6 91.1 65.2 71.8 86.0 25.1 67.4 46.0 51.9 69.1 66.82 62.7 49.4 50.3 78.0 60.09
Table 8: Experiments using a different source dataset (a 16-shot subset of LAION-2B queried using ImageNet label names). Settings are identical to Table 10 (the expanded form of Table 6 in the main paper).
@@ Appendix C Token offset trick (for Descriptor Soup)
Figure 7: Different number of shots. We experiment with the same 14 datasets as the main paper and report average of 3 random trials. We report average target accuracies over 10 diverse datasets (left) and 4 ImageNet shifts (right). Here we verify that the improvements of both word soup and CoOp + word soup over CoOp are resilient to the number of shots. Indeed, we emphasize that word soup is very resilient in extreme low shot scenarios due to the low number of parameters.
Appendix E Additional Ablation Studies
We present additional ablation studies in Table 8 and Figure 7. Table 8 presents OOD generalization results with a different source data set. Figure 7 presents results with different number of shots.
Parameter Efficiency
Computational Efficiency
@@ Appendix E Additional Ablation Studies
CoOp [70$\dagger$ 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88 64.20 47.99 49.71 75.21 59.3
Co-CoOp [69$\dagger$ 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74 64.07 48.75 50.63 76.18 59.9